# Exponential Stability of Parametric Optimization-Based Controllers via Lur’e Contractivity

### v1

@@ I Introduction
As our second main contribution, we establish in Theorem 1 a novel necessary and sufficient condition for contractivity of Lur’e systems with monotone and Lipschitz nonlinearities. Our condition improves upon the condition in [2] and [10, Proposition 4], whereby we demonstrate that their condition is a special case of ours. Additionally, our result is stronger than [7, Theorem 4.2] since our condition is necessary and sufficient.
@@ II-B Virtual System Method for Convergence Analysis
For example, if $d=n$ and $f(x)=f_{\textup{virtual}}(x,x)$, then one can see that $\phi_{t}(x_{0})$ is a solution for both systems and is often selected as one of the two specific solutions in step (ii). Additionally, if $f_{\textup{virtual}}(\mathbbold{0}_{n},z)=\mathbbold{0}_{n}$ for all $z\in{}^{n}$, then $\mathbbold{0}_{n}$ is an equilibrium point for the virtual system and can be selected as one of the specific solutions.
@@ III Absolute Contractivity of Lur’e Systems
where $\varphi:{}^{m}\times\mathbb{R}_{\geq 0}\rightarrow{}^{m}$ is Lipschitz in its first argument and continuous in its second argument. We additionally assume that $\varphi$ is monotone in its first argument, i.e., there exists a constant $\rho>0$ such that for all $y_{1},y_{2}\in{}^{m},t\geq 0$,
$0\leq(\varphi(y_{1},t)-\varphi(y_{2},t))^{\top}(y_{1}-y_{2})\leq\rho\|y_{1}-y_{2}\|_{2}^{2}.$ (6)
@@
$\begin{bmatrix}A^{\top}P+PA+2\eta P&PB+\lambda K^{\top}\\ B^{\top}P+\lambda K&-2\lambda\rho^{-1}I_{m}\end{bmatrix}\preceq 0.$ (7)
@@
    $\Delta x^{\top}(PA+A^{\top}P+2cP)\Delta x+2\Delta x^{\top}PB\Delta u_{t}\leq 0.$ (8)
    Moreover, we claim that the monotonicity condition (6) is equivalent to
    $\Delta u_{t}^{\top}(\rho^{-1}\Delta u_{t}-K\Delta x)\leq 0$ (9)
    $\quad\iff\qquad\rho^{-1}\|\Delta u_{t}\|_{2}^{2}\leq\Delta u_{t}^{\top}\Delta y$ (10)
    To see this fact, note that the inequality (10) is trivially true when $\Delta u_{t}=\mathbbold{0}_{m}$, so we assume $\Delta u_{t}\neq\mathbbold{0}_{m}$. Inequality (10) directly implies the left inequality of (6). Additionally, multiplying both sides of (10) by $\rho\Delta u_{t}^{\top}\Delta y/\|\Delta u_{t}\|_{2}^{2}>0$ implies the right inequality of (6) by Cauchy-Schwarz. The reverse implication holds by multiplying the inequality (6) by the positive quantity $\rho^{-1}\Delta u_{t}^{\top}\Delta y/\|\Delta y\|_{2}^{2}$ (noting that the result is trivial when $\Delta y=\mathbbold{0}_{n}$ since $\Delta y=\mathbbold{0}_{m}\implies\Delta u_{t}=\mathbbold{0}_{m}$). Asking when the inequality (9) implies (8) is equivalent to the inequality (7) in light of the necessity and sufficiency of the S-procedure [17]. ∎
Note that the condition in [2, Theorem 2] corresponds to the inequality (7) with $\lambda=1$. Moreover the matrix in (7) has $A^{\top}P+PA+2\eta P$ in its $(1,1)$ block compared to $A^{\top}P+PA+\eta I_{n}$ in [2]. This modification ensures that the inequality (1) holds rather than a related inequality with $-\frac{\eta}{2}\|x_{1}-x_{2}\|_{2}^{2}$ on the right-hand side.
@@ IV-B Stability Analysis for LTI Systems
    Note that for all $t\in[0,\tau_{\textup{max}}(x_{0}))$, $\operatorname{Proj}_{\Gamma(\phi_{x_{0}}(t))}$ obeys the inequality (6) with $\rho=1$ due to monotonicity and nonexpansiveness, see, e.g., [20, Eq. (2)]. Therefore the virtual system is a Lur’e system of the form (5). Theorem 1 implies{arxiv}  [^2] To apply Theorem 1, we need to establish Lipschitzness in $y$ and continuity in $t$ for the virtual system (17). We prove these properties in Lemma 16. that this virtual system is contracting with respect to $\|\cdot\|_{P}$ with rate $\eta>0$. In other words, any two trajectories $y_{1}(\cdot),y_{2}(\cdot)$ for the virtual system satisfy for all $t\in[0,\tau_{\textup{max}}(x_{0}))$,
@@ V-B Stability with Control Barrier Functions
A continuous controller $u:{}^{n}\rightarrow{}^{m}$ which satisfies (23) for all $x\in\mathcal{C}$ renders $\mathcal{C}$ forward-invariant under the dynamics (22[13, Theorem 4].
@@
Example 8 ():
Consider a two-dimensional single integrator attempting to avoid a disk-shaped obstacle centered at $(0,4)$ with radius $2$. The corresponding CBF is $h(x_{1},x_{2})=x_{1}^{2}+(x_{2}-4)^{2}-4$ with $\alpha(r)=r$. We take $K=\left[\begin{smallmatrix}-2&-0.5\\ -0.5&-1\end{smallmatrix}\right]$ and $\bar{u}=1$. We plot numerical simulations of (24) with these parameters along with the corresponding convergence rate and value of CBF along trajectories in Figure 2. We can observe that all trajectories converge to the set of equilibria and the majority of them converge to the origin. Moreover, we see that the exponential convergence rate predicted in (25) appears to be tight in this example and that all trajectories remain in $\mathrm{int}(\mathcal{C})$ for all $t\geq 0$. While the equilibrium point on the boundary of the safe set is unavoidable in this example, we can numerically observe that there are no equilibria inside $\operatorname{int}(\mathcal{C})$ other than the origin. This result is in contrast with the CLF-CBF controllers studied in [19, 21], whereby there may exist additional undesirable equilibria inside $\operatorname{int}(\mathcal{C})$. This numerical example provides evidence that one does not need to enforce any CLF decrease condition to ensure convergence to the origin, except on a set of measure zero.
@@ VI Discussion and Future Work
We recall some definitions from set-valued analysis and refer to [8, Section 2.1.3] for a more comprehensive treatment. A point-to-set map, also called a set-valued map, is a map, $K$, from ${}^{n}$ to the power set of ${}^{m}$, i.e., the set of subsets of ${}^{m}$. We will denote a point-to-set map by $K:{}^{n}\rightrightarrows{}^{m}$. The domain of $K$ is denoted $\operatorname{Dom}(K)$ and is the set $\operatorname{Dom}(K):=\{x\in{}^{n}\;|\;K(x)\neq\emptyset\}$.
@@
Definition 10 (Lower semicontinuity):
@@
We say a set-valued map $K:{}^{n}\rightrightarrows{}^{m}$ is continuous at $x$ if it is both lower semicontinuous and upper semicontinuous at $x$. Suppose that $K:{}^{n}\rightrightarrows{}^{m}$ is a closed-valued set-valued map, i.e., $K(x)$ is a closed set for all $x\in\operatorname{Dom}(K)$. Then for every $x\in\operatorname{Dom}(K)$, the distance function $\mathrm{dist}(\cdot,K(x))$ is well-defined, where
@@
Note that for every $x\in{}^{n}$, $\liminf_{y\to x}K(y)\subseteq K(x)\subseteq\limsup_{y\to x}K(y)$. If equality holds between the liminf and limsup sets, we write the common set as $\lim_{y\to x}K(y)$ and call it the limit of $K(y)$ as $y$ tends to $x$. We will now require some auxiliary results.
@@
    We simply need to extend Proposition 13 to the case that $g$ is vector-valued. For $i\in\{1,\dots,p\}$, define the set-valued maps $\Gamma_{i}(x):=\{u\in{}^{m}\;|\;g_{i}(x,u)\leq 0\}$ and note that $\Gamma(x)=\bigcap_{i=1}^{p}\Gamma_{i}(x)$. We simply need to show upper and lower semicontinuity of $\Gamma$ to conclude continuity. Regarding lower-semicontinuity, let $\bar{x}\in\operatorname{Dom}(\Gamma)$ be such that there exists $\bar{u}$ with $g(\bar{x},\bar{u})\ll\mathbbold{0}_{p}$ and let $\mathcal{X}\subseteq{}^{n}$ be open and satisfy $\Gamma(\bar{x})\cap\mathcal{X}\neq\emptyset$. By continuity of each of the $\Gamma_{i}$ at $\bar{x}$ (due to Proposition 13), there exists an open set $\mathcal{N}_{i}$ containing $\bar{x}$ such that for each $x^{i}\in\mathcal{N}_{i}$, $\Gamma_{i}(x^{i})\cap\mathcal{X}\neq\emptyset$ for all $i\in\{1,\dots,p\}$. Since a finite intersection of open sets is open, let $\mathcal{N}:=\bigcap_{i=1}^{p}\mathcal{N}_{i}$. Then it is routine to check that for each $x\in\mathcal{N}$, $\Gamma(x)\cap\mathcal{X}\neq\emptyset$. This establishes lower semicontinuity. To establish upper semicontinuity, let $\bar{x}$ satisfy the stated assumptions and let $\mathcal{V}$ be an open set with $\Gamma(\bar{x})\subseteq\mathcal{V}$. By continuity of each of the $\Gamma_{i}$ at $\bar{x}$, there exists an open neighborhood $\mathcal{N}_{i}$ of $\bar{x}$ such that for each $x^{i}\in\mathcal{N}_{i}$, $\Gamma_{i}(x)\subseteq\mathcal{V}$. Once again taking $\mathcal{N}=\bigcap_{i=1}^{p}\mathcal{N}$ ensures that for all $x\in\mathcal{N}$, we have that $\Gamma(x)\subseteq\mathcal{V}$. This establishes upper semicontinuity and thus continuity of $\Gamma$ at $\bar{x}$. The result is then a consequence of Propositions 11 and 12. ∎
@@
Then for each $a\in\mathcal{A}$, there exists a unique $x^{\star}(a)$ satisfying the fixed point equation $x^{\star}(a)=f(x^{\star}(a),a)$ and the mapping $a\mapsto x^{\star}(a)$ is continuous. Moreover, if there exists $L>0$ such that for each $x\in{}^{n}$, the map $a\mapsto f(x,a)$ is Lipschitz with constant $L$, then the mapping $a\mapsto x^{\star}(a)$ is Lipschitz with estimate $L/(1-\ell)$.
@@
    where $\gamma>0$ is a constant to be tuned to make the map contractive. Let $\bar{f}:{}^{m}\times{}^{n}\times[0,\tau_{\textup{max}}(x_{0}))\rightarrow{}^{m}$ denote this fixed-point map, i.e., $\bar{f}(u,z,t):=\operatorname{Proj}_{\Gamma(\phi_{x_{0}}(t))}(u-\gamma\nabla_{u}f(z,u)).$ Under the assumption that $\nabla_{u}f$ is Lipschitz, from standard results on projected gradient algorithms, e.g., [20, pp. 31], we know that for fixed $z,t$, the map $u\mapsto\bar{f}(u,z,t)$ is a contraction for $0<\gamma<2/L$, where $L$ is the Lipschitz constant of $\nabla_{u}f(z,\cdot)$. Henceforth, we let $\gamma$ be in this range. To establish Lipschitzness in $z$, we note that at fixed $u,t$, the mapping $z\mapsto\bar{f}(u,z,t)$ is Lipschitz under the assumption that the mapping $\nabla_{u}f(\cdot,u)$ is Lipschitz. Thus, at fixed $t$, we can apply Proposition 15 to conclude that the mapping $z\mapsto\bar{u}(z,t)$ is Lipschitz and thus continuous in its first argument. To establish continuity in $t$, we first study the point-to-set mapping $t\rightrightarrows\Gamma(\phi_{x_{0}}(t))$. By assumption that for all $t$, we have that $\phi_{x_{0}}(t)\in\mathcal{K}$, we conclude by Lemma 14 that the mapping $\Phi:{}^{m}\times[0,\tau_{\textup{max}}(x_{0}))\rightarrow{}^{m}$ given by $\Phi(u,t):=\operatorname{Proj}_{\Gamma(\phi_{x_{0}}(t))}(u)$ is continuous on its domain. Since $\bar{f}(u,z,t)=\Phi(u-\gamma\nabla_{u}f(z,u),t)$, we have that $\bar{f}(u,z,\cdot)$ is continuous as well. And since for all $z,t$, we know that $u\mapsto\bar{f}(u,z,t)$ is a contraction, by Proposition 15, we know that for all $z\in{}^{m}$, the map $t\mapsto\bar{u}(z,t)$ is continuous. ∎

### v2

@@ I Introduction
As our second main contribution, we establish in Theorem 1 a novel necessary and sufficient condition for contractivity of Lur’e systems with cocoercive nonlinearities. In contrast, in [2] and [10, Proposition 4], monotone and Lipschitz nonlinearities are considered yet only sufficient conditions are provided. In other words, by studying a smaller class of nonlinearities, we are able to find the sharpest condition for contractivity. Additionally, our result is stronger than [7, Theorem 4.2] since our condition is necessary and sufficient.
@@ II-B Virtual System Method for Convergence Analysis
For example, if $d=n$ and $f(x)=f_{\textup{virtual}}(x,x)$, then one can see that $\phi_{x_{0}}(t)$ is a solution for both systems and is often selected as one of the two specific solutions in step (ii). Additionally, if $f_{\textup{virtual}}(\mathbbold{0}_{n},z)=\mathbbold{0}_{n}$ for all $z\in{}^{n}$, then $\mathbbold{0}_{n}$ is an equilibrium point for the virtual system and can be selected as one of the specific solutions.
@@ III Absolute Contractivity of Lur’e Systems
where $\varphi:{}^{m}\times\mathbb{R}_{\geq 0}\rightarrow{}^{m}$ is cocoercive in its first argument and continuous in its second argument. Specifically, there exists a constant $\rho>0$ such that for all $y_{1},y_{2}\in{}^{m},t\geq 0$,
$(\varphi(y_{1},t)-\varphi(y_{2},t))^{\top}(y_{1}-y_{2})\geq\rho\|\varphi(y_{1},t)-\varphi(y_{2},t)\|_{2}^{2}.$ (6)
Notably, cocoercivity, (6), implies that $\varphi$ is monotone and Lipschitz continuous with constant $\rho^{-1}$ in its first entry. Many standard nonlinearities satisfy cocoercivity including projections onto convex sets and nonlinearities of the form $\varphi(x)=(\varphi_{1}(x_{1}),\dots,\varphi_{m}(x_{m}))$ where each $\varphi_{i}$ is slope-restricted between 00 and $\rho^{-1}$.
@@
$\begin{bmatrix}A^{\top}P+PA+2\eta P&PB+\lambda K^{\top}\\ B^{\top}P+\lambda K&-2\lambda\rho I_{m}\end{bmatrix}\preceq 0.$ (7)
@@
    $\Delta x^{\top}(PA+A^{\top}P+2\eta P)\Delta x+2\Delta x^{\top}PB\Delta u_{t}\leq 0.$ (8)
    Moreover, the cocoercivity condition (6) is equivalent to
    $\Delta u_{t}^{\top}(\rho\Delta u_{t}-K\Delta x)\leq 0.$ (9)
    Asking when the inequality (9) implies (8) is equivalent to the inequality (7) in light of the necessity and sufficiency of the S-procedure [17]. ∎
Note that the condition in [2, Theorem 2] corresponds to the inequality (7) with $\lambda=1$. Moreover the matrix in (7) has $A^{\top}P+PA+2\eta P$ in its $(1,1)$ block compared to $A^{\top}P+PA+\eta I_{n}$ in [2]. This modification ensures that the inequality (1) holds rather than a related inequality with $-\frac{\eta}{2}\|x_{1}-x_{2}\|_{2}^{2}$ on the right-hand side. On the other hand, there exist mappings $\varphi$ that are monotone and Lipschitz but not cocoercive. Thus, we consider a smaller class of mappings in Theorem 1, but are able to provide necessary and sufficient conditions for absolute contractivity.
@@ IV-B Stability Analysis for LTI Systems
    Note that for all $t\in[0,\tau_{\textup{max}}(x_{0}))$, $\operatorname{Proj}_{\Gamma(\phi_{x_{0}}(t))}$ obeys the inequality (6) with $\rho=1$ due to cocoercivity of projections, see, e.g., [20, Eq. (2)]. Therefore the virtual system is a Lur’e system of the form (5). Theorem 1 implies{arxiv}  [^2] To apply Theorem 1, we need to establish continuity in $t$ for the virtual system (16). We prove this property in Lemma 16. that this virtual system is contracting with respect to $\|\cdot\|_{P}$ with rate $\eta>0$. In other words, any two trajectories $y_{1}(\cdot),y_{2}(\cdot)$ for the virtual system satisfy for all $t\in[0,\tau_{\textup{max}}(x_{0}))$,
@@ V-A State-Dependent Saturation Control
We can see that all trajectories converge exponentially quickly to the origin and that the estimated exponential convergence rate from Theorem 3 is $\eta=0.0525$. Empirically, we can see that when trajectories are far from the origin, this rate is tight since $v(x)\approx\mathbbold{0}_{m}$ for $x$ far from $\mathbbold{0}_{3}$ and that the rate is very loose when trajectories are close to the origin since $v(x)\approx\mathbbold{1}_{m}$ for $x\approx\mathbbold{0}_{3}$.
@@ V-B Stability with Control Barrier Functions
A continuous controller $u:{}^{n}\rightarrow{}^{m}$ which satisfies (22) for all $x\in\mathcal{C}$ renders $\mathcal{C}$ forward-invariant under the dynamics (21[13, Theorem 4].
@@
Example 8:
Consider a two-dimensional single integrator attempting to avoid a disk-shaped obstacle centered at $(0,4)$ with radius $2$. The corresponding CBF is $h(x_{1},x_{2})=x_{1}^{2}+(x_{2}-4)^{2}-4$ with $\alpha(r)=r$. We take $K=\left[\begin{smallmatrix}-2&-0.5\\ -0.5&-1\end{smallmatrix}\right]$ and $\bar{u}=1$. We plot numerical simulations of (23) with these parameters along with the corresponding convergence rate and value of CBF along trajectories in Figure 2.
We can observe that all trajectories converge to the set of equilibria and the majority of them converge to the origin. Moreover, we see that the exponential convergence rate predicted in (24) appears to be tight in this example and that all trajectories remain in $\mathrm{int}(\mathcal{C})$ for all $t\geq 0$. While the equilibrium point on the boundary of the safe set is unavoidable in this example, we can numerically observe that there are no equilibria inside $\operatorname{int}(\mathcal{C})$ other than the origin. This result is in contrast with the CLF-CBF controllers studied in [19, 21], whereby there may exist additional undesirable equilibria inside $\operatorname{int}(\mathcal{C})$. This numerical example provides evidence that one does not need to enforce any CLF decrease condition to ensure convergence to the origin, except on a set of measure zero.
@@ VI Discussion and Future Work
We recall some definitions from set-valued analysis and refer to [8, Section 2.1.3] for a more comprehensive treatment.
A point-to-set map, also called a set-valued map, is a map, $K$, from n to the power set of m, i.e., the set of subsets of m. We will denote a point-to-set map by $K:{}^{n}\rightrightarrows{}^{m}$. The domain of $K$ is denoted $\operatorname{Dom}(K)$ and is the set $\operatorname{Dom}(K):=\{x\in{}^{n}\;|\;K(x)\neq\emptyset\}$.
@@
Definition 10 (Upper semicontinuity):
@@
We say a set-valued map $K:{}^{n}\rightrightarrows{}^{m}$ is continuous at $x$ if it is both lower semicontinuous and upper semicontinuous at $x$.
Suppose that $K:{}^{n}\rightrightarrows{}^{m}$ is a closed-valued set-valued map, i.e., $K(x)$ is a closed set for all $x\in\operatorname{Dom}(K)$. Then for every $x\in\operatorname{Dom}(K)$, the distance function $\mathrm{dist}(\cdot,K(x))$ is well-defined, where
@@
Note that for every $x\in{}^{n}$, $\liminf_{y\to x}K(y)\subseteq K(x)\subseteq\limsup_{y\to x}K(y)$. If equality holds between the liminf and limsup sets, we write the common set as $\lim_{y\to x}K(y)$ and call it the limit of $K(y)$ as $y$ tends to $x$.
We will now require some auxiliary results.
@@
    We simply need to extend Proposition 13 to the case that $g$ is vector-valued. For $i\in\{1,\dots,p\}$, define the set-valued maps $\Gamma_{i}(x):=\{u\in{}^{m}\;|\;g_{i}(x,u)\leq 0\}$ and note that $\Gamma(x)=\bigcap_{i=1}^{p}\Gamma_{i}(x)$. We simply need to show upper and lower semicontinuity of $\Gamma$ to conclude continuity. Regarding lower-semicontinuity, let $\bar{x}\in\operatorname{Dom}(\Gamma)$ be such that there exists $\bar{u}$ with $g(\bar{x},\bar{u})\ll\mathbbold{0}_{p}$ and let $\mathcal{X}\subseteq{}^{n}$ be open and satisfy $\Gamma(\bar{x})\cap\mathcal{X}\neq\emptyset$. By continuity of each of the $\Gamma_{i}$ at $\bar{x}$ (due to Proposition 13), there exists an open set $\mathcal{N}_{i}$ containing $\bar{x}$ such that for each $x^{i}\in\mathcal{N}_{i}$, $\Gamma_{i}(x^{i})\cap\mathcal{X}\neq\emptyset$ for all $i\in\{1,\dots,p\}$. Since a finite intersection of open sets is open, let $\mathcal{N}:=\bigcap_{i=1}^{p}\mathcal{N}_{i}$. Then it is routine to check that for each $x\in\mathcal{N}$, $\Gamma(x)\cap\mathcal{X}\neq\emptyset$. This establishes lower semicontinuity.
    To establish upper semicontinuity, let $\bar{x}$ satisfy the stated assumptions and let $\mathcal{V}$ be an open set with $\Gamma(\bar{x})\subseteq\mathcal{V}$. By continuity of each of the $\Gamma_{i}$ at $\bar{x}$, there exists an open neighborhood $\mathcal{N}_{i}$ of $\bar{x}$ such that for each $x^{i}\in\mathcal{N}_{i}$, $\Gamma_{i}(x)\subseteq\mathcal{V}$. Once again taking $\mathcal{N}=\bigcap_{i=1}^{p}\mathcal{N}$ ensures that for all $x\in\mathcal{N}$, we have that $\Gamma(x)\subseteq\mathcal{V}$. This establishes upper semicontinuity and thus continuity of $\Gamma$ at $\bar{x}$.
    The result is then a consequence of Propositions 11 and 12. ∎
@@
Then for each $a\in\mathcal{A}$, there exists a unique $x^{\star}(a)$ satisfying the fixed point equation $x^{\star}(a)=f(x^{\star}(a),a)$ and the mapping $a\mapsto x^{\star}(a)$ is continuous.
Moreover, if there exists $L>0$ such that for each $x\in{}^{n}$, the map $a\mapsto f(x,a)$ is Lipschitz with constant $L$, then the mapping $a\mapsto x^{\star}(a)$ is Lipschitz with estimate $L/(1-\ell)$.
@@
    where $\gamma>0$ is a constant to be tuned to make the map contractive. Let $\bar{f}:{}^{m}\times{}^{n}\times[0,\tau_{\textup{max}}(x_{0}))\rightarrow{}^{m}$ denote this fixed-point map, i.e., $\bar{f}(u,z,t):=\operatorname{Proj}_{\Gamma(\phi_{x_{0}}(t))}(u-\gamma\nabla_{u}f(z,u)).$ Under the assumption that $\nabla_{u}f$ is Lipschitz, from standard results on projected gradient algorithms, e.g., [20, pp. 31], we know that for fixed $z,t$, the map $u\mapsto\bar{f}(u,z,t)$ is a contraction for $0<\gamma<2/L$, where $L$ is the Lipschitz constant of $\nabla_{u}f(z,\cdot)$. Henceforth, we let $\gamma$ be in this range. To establish Lipschitzness in $z$, we note that at fixed $u,t$, the mapping $z\mapsto\bar{f}(u,z,t)$ is Lipschitz under the assumption that the mapping $\nabla_{u}f(\cdot,u)$ is Lipschitz. Thus, at fixed $t$, we can apply Proposition 15 to conclude that the mapping $z\mapsto\bar{u}(z,t)$ is Lipschitz and thus continuous in its first argument.
    To establish continuity in $t$, we first study the point-to-set mapping $t\rightrightarrows\Gamma(\phi_{x_{0}}(t))$. By assumption that for all $t$, we have that $\phi_{x_{0}}(t)\in\mathcal{K}$, we conclude by Lemma 14 that the mapping $\Phi:{}^{m}\times[0,\tau_{\textup{max}}(x_{0}))\rightarrow{}^{m}$ given by $\Phi(u,t):=\operatorname{Proj}_{\Gamma(\phi_{x_{0}}(t))}(u)$ is continuous on its domain. Since $\bar{f}(u,z,t)=\Phi(u-\gamma\nabla_{u}f(z,u),t)$, we have that $\bar{f}(u,z,\cdot)$ is continuous as well. And since for all $z,t$, we know that $u\mapsto\bar{f}(u,z,t)$ is a contraction, by Proposition 15, we know that for all $z\in{}^{m}$, the map $t\mapsto\bar{u}(z,t)$ is continuous. ∎